### v1

@@ 7.1 Isolated and Continuous SLR
Name Task Year Modality Samples Signers Sign Voc. Language Duration Source
PHOENIX-2014 [44] CSLR/SLT 2014 RGB - 9 1558 DGS 10.73h TV
PHOENIX-2014-T1 [26] CSLR/SLT 2018 RGB 8257 9 1066 DGS - TV
KETI [90] CSLR/SLT 2019 RGB, Pose 14672 14 524 KSL 27.99h Lab
GSL SD/SI [2] CSLR/SLT 2019 RGBD 10295 7 310 GSL 9.59h Lab
Corpus FinSL [154] CSLR/SLT 2020 RGB 343 21 - FinSL 14.37h -
CSL-Daily [209] CSLR/SLT 2021 RGB 20654 10 2000 CSL 23.27h Lab
How2Sign (green studio) [41] CSLR/SLT 2021 RGBD, Pose 2529 11 - ASL 79.12h Lab
BOBSL [12] CSLR/SLT 2021 RGB 1962 39 2281 BSL 1467h TV
OpenASL [160] CSLR/SLT 2022 RGB 98417 220 - ASL 288h Web
Auslan-Daily [158] CSLR/SLT 2023 RGB 25K 67 - AUSLAN - TV, Web
TVB-HKSL-News [130] CSLR/SLT 2024 RGB, Pose 7K 2 6515 HKSL 16.07h TV
QSL [48] CSLR/SLT 2024 RGB 30040 5 654 CSL - Lab
ChicagoFSWild+ [159] CSLR 2019 RGB 55232 260 26 ASL - Web
HKSL [212] CSLR 2021 RGBD 2400 6 55 HKSL - Lab
CSL [78] IDSLR 2018 RGBD, Pose 125K 50 500 CSL - Lab
SMILE [42] IDSLR 2018 RGB - 30 100 DSGS - Lab
ASL Dataset [16] IDSLR 2018 Pose 1200 20 30 ASL - Lab
MSASL [85] IDSLR 2019 RGB 25513 222 1000 ASL 24.65h Web
WLASL [99] IDSLR 2019 RGB 21083 119 2000 ASL 14.11h Web
GSL isol. [2] IDSLR 2019 RGBD 40785 7 310 GSL 6.44h Lab
BosphorusSign22k [217] IDSLR 2020 RGBD, Pose 22542 6 744 TSL 19h Lab
INCLUDE [169] IDSLR 2020 RGB 4287 7 263 ISL 3h Lab
AUTSL [164] IDSLR 2020 RGBD, Pose 38336 43 226 TSL 21h Lab
RKS-PERSIANSIGN [144] IDSLR 2020 RGB 10000 10 100 Persian - Lab
BSL-1K [11] IDSLR 2020 RGB 280K 40 1064 BSL 1060h TV
NMFs-CSL [69] IDSLR 2021 RGB 32010 10 1.067 CSL 24.4h Lab
KArSL [163] IDSLR 2021 RGBD, Pose 75300 3 502 ArSL - Lab
CSLD [184] IDSLR 2022 RGB 120K 30 400 CSL 66.8h Lab
NCSL [182] IDSLR 2022 RGB 90K 30 300 CSL 50.1h Lab
IISL2020 [92] IDSLR 2022 RGB 12100 16 11 ISL 7h Lab
LSA64 [149] IDSLR 2023 RGB 3200 10 64 LSA 1.9h Lab
@@
Signsuisse Lexicon [126] IDSLR 2023 RGB 18221 - - 3 (Mult) - Web
FDMSE-ISL [135] IDSLR 2024 RGBD 40033 20 2002 ISL 36h Lab
@@ 7.2 Semaphoric HGR
Semaphoric hand gesture recognition (HGR) research is supported by a diverse range of datasets categorized into three main types based on task formulation: static, isolated dynamic, and continuous dynamic gesture recognition. These datasets vary in their application domains (e.g., Human-Computer Interaction (HCI), Human-Machine Interface (HMI), Human-Robot Interaction (HRI)), modalities (e.g., RGB, depth, pose, infrared), and in the number of gesture classes, subjects, and gesture samples they include. The statistics of datasets for semaphoric HGR are provided in tables 9,10 and 11, regarding the task for which they are used (i.e. static, isolated dynamic and continuous HGR, respectively).
Static HGR datasets focus on recognizing gestures from single-frame input and are generally easier to process due to the absence of temporal complexity. Notable examples include LaRED [67], one of the earliest RGBD datasets with 243,000 samples and 81 classes, and HaGRIDv2 [131], a large-scale RGB dataset designed for HMI with over 1 million samples, covering 34 classes. Smaller yet targeted datasets such as OUHANDS [118] and URGR [17] focus on specific domains like HCI and HRI, while MD-UHGRD [194] offers moderately sized samples and class diversity. Isolated dynamic HGR datasets contain gestures performed in isolation, with start and end points explicitly segmented. Datasets like ChaLearn IsoGD [179] and EgoGesture (isolated) [199] feature rich modality (RGBD, IR, pose) and high gesture diversity, with up to 249 classes. Others like Jester [117] offer large-scale RGB-only data from over 1,300 subjects, making them suitable for data-driven deep learning approaches. More recent contributions such as DATE [35] and ZJUGesture [193] reflect growing interest in real-world HCI scenarios. Continuous dynamic HGR datasets challenge models to identify and segment gestures from continuous video streams. The ChaLearn ConGD [179] dataset remains a benchmark with 22,535 sequences and 249 gesture classes. Similarly, EgoGesture [199] and IPN Hand [19] include both isolated and continuous segments, enabling comprehensive model development and evaluation. Datasets like LD-ConGR [106] and ZJUGesture [193] contribute additional complexity by increasing subject diversity and contextual variability.
@@
Table 9: Datasets for static hand gesture recognition.
Name Year Domain Modality Samples Subjects Classes Resolution
HGR-1 [88] 2014 - RGB 899 12 27 174x131-640x480
LaRED [67] 2014 - RGBD 243000 10 81 640x480
OUHANDS [118] 2016 HCI RGBD 3150 23 10 640x480
LRHG [211] 2021 HRI RGB 4320 8 10 640x480
SHAPE [33] 2022 - RGB 34226 20 32 4128x3096
HaGRID [86] 2024 HCI RGB 552992 34730 19 1920x1080
HaGRIDv2 [131] 2024 HCI RGB 1086158 65977 34 1920x1080
MD-UHGRD [194] 2024 HRI RGB 20000 - 19 -
URGR [17] 2024 HRI RGB 347483 16 6 480x640
Table 10: Datasets for isolated dynamic hand gesture recognition.
@@
SKIG [108] 2013 - RGBD 1080 6 10 -
nvGesture [124] 2016 HMI RGBD, IR, OF 1532 20 25 -
ChaLearn IsoGD [179] 2016 - RGBD 47933 21 249 -
DHG-14/28 [37] 2016 - Depth, Pose 2800 20 14/28 -
SHREC17-14/28 [166] 2017 - Depth, Pose 2800 28 14/28 -
EgoGesture (isol.) [199] 2017 HCI RGBD 24161 50 83 2953224
Briareo [115] 2019 HMI RGBD, Pose, IR 1440 40 12 -
Jester [117] 2019 HCI RGB 148092 1376 27 5331312
IPN Hand (isol.) [19] 2020 HCI RGB 4218 50 13 -
TCG [189] 2020 HMI Pose 250 5 4 839350
LD-ConGR (isol.) [106] 2022 HCI RGBD 44887 30 10 -
ZJUGesture (isol.) [193] 2023 HCI RGB 9892 60 9 -
DATE [35] 2024 HCI RGB 13500 22 27 -
EgoDriving [138] 2024 HMI RGB 400 - 29 70276
Table 11: Datasets for continuous dynamic hand gesture recognition.
@@
ChaLearn ConGD [179] 2016 - RGBD 22535 47933 21 249
EgoGesture (cont.) [199] 2017 HCI RGBD 2081 24161 50 83
IPN Hand (cont.) [19] 2020 HCI RGB 200 4218 50 13
Police Gesture Dataset [66] 2020 HMI RGB 21 3354 - 8
LD-ConGR (cont.) [106] 2022 HCI RGBD 542 44887 30 10
ZJUGesture (cont.) [193] 2023 HCI RGB - 9892 60 9
@@ 7.3 Other
In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types. The First-Person Hand Action (FPHA) dataset [50], published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants. It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction. Another notable example is the DP Dataset [127], released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition. Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances. It includes detailed annotations such as gesture timings and pointing directions, enabling both binary gesture classification (gesture vs. non-gesture) and directional estimation tasks.

### v2

@@ 8.1 Isolated and Continuous SLR
Name Year Task Modality Samples Signers Sign Voc. Language Duration Source
TVB-HKSL-News (TVB-HKSL-News) 2024 CSLR/SLT RGB, Pose 7000 2 6515 HKSL 16.07h TV
Auslan-Daily (Auslan-Daily) 2023 CSLR/SLT RGB 25000 67 - AUSLAN - TV, Web
OpenASL (OpenASL) 2022 CSLR/SLT RGB 98417 220 - ASL 288h Web
CSL-Daily (Back-Translation-CVPR) 2021 CSLR/SLT RGB 20654 10 2000 CSL 23.27h Lab
BOBSL (BOBSL-BBC) 2021 CSLR/SLT RGB 1962 39 2281 BSL 1467h TV
How2Sign (green studio) (How2Sign-CVPR) 2021 CSLR/SLT RGBD, Pose 2529 11 - ASL 79.12h Lab
Corpus FinSL (Corpus-Finnish) 2020 CSLR/SLT RGB 343 21 - FinSL 14.37h -
KETI (KETI_MDPI) 2019 CSLR/SLT RGB, Pose 14672 14 524 KSL 27.99h Lab
GSL SD/SI (ADADOGLOU) 2019 CSLR/SLT RGBD 10295 7 310 GSL 9.59h Lab
PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) 2018 CSLR/SLT RGB 8257 9 1066 DGS - TV
PHOENIX-2014 (PHOENIX-2014) 2014 CSLR/SLT RGB - 9 1558 DGS 10.73h TV
HKSL (SignBERT-HKSL) 2021 CSLR RGBD 2400 6 55 HKSL - Lab
ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) 2019 CSLR RGB 55232 260 26 ASL - Web
Multi-VSL (Multi-VSL-dinh2025sign) 2025 IDSLR RGB, Pose 84764 30 1000 Vietnamese - Lab
BdSLW60 (BdSLW60-rubaiyeat2025bdslw60) 2025 IDSLR RGB 9307 18 60 Bangla - Lab
ISLR101 (ISLR101-IRANIAN) 2025 IDSLR RGB, Pose 4614 10 101 Iranian - -
ISW-1000 (ISW-1000-ren2025multi) 2025 IDSLR RGB, Pose, OF 10000 10 1000 CSL - Lab
FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention) 2024 IDSLR RGBD 40033 20 2002 ISL 36h Lab
LSA64 (Lsa64argentiniansignlanguage) 2023 IDSLR RGB 3200 10 64 LSA 1.9h Lab
@@
Signsuisse Lexicon (SIGN-muller-etal-2023-findings) 2023 IDSLR RGB 18221 - - 3 (Mult) - Web
CSLD (tim-slr_csld_dataseta) 2022 IDSLR RGB 120000 30 400 CSL 66.8h Lab
NCSL (2plus1-D-SLR_and_NCSL) 2022 IDSLR RGB 90000 30 300 CSL 50.1h Lab
KArSL (KArSL) 2021 IDSLR RGBD, Pose 75300 3 502 ArSL - Lab
NMFs-CSL (NMFs-CSL) 2021 IDSLR RGB 32010 10 1.067 CSL 24.4h Lab
BosphorusSign22k (bosphorussign22k) 2020 IDSLR RGBD, Pose 22542 6 744 TSL 19h Lab
AUTSL (AUTSL) 2020 IDSLR RGBD, Pose 38336 43 226 TSL 21h Lab
INCLUDE (INCLUDE) 2020 IDSLR RGB 4287 7 263 ISL 3h Lab
IISL2020 (IISL2020_MDPI) 2020 IDSLR RGB 12100 16 11 ISL 7h Lab
RKS-PERSIANSIGN (persian_sl_dataset) 2020 IDSLR RGB 10000 10 100 Persian - Lab
BSL-1K (bsl-1K) 2020 IDSLR RGB 280000 40 1064 BSL 1060h TV
MSASL (MS-ASL) 2019 IDSLR RGB 25513 222 1000 ASL 24.65h Web
WLASL (WLASL) 2019 IDSLR RGB 21083 119 2000 ASL 14.11h Web
GSL isol. (ADADOGLOU) 2019 IDSLR RGBD 40785 7 310 GSL 6.44h Lab
SMILE (SMILE) 2018 IDSLR RGB - 30 100 DSGS - Lab
ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) 2018 IDSLR Pose 1200 20 30 ASL - Lab
CSL (CSL500) 2018 IDSLR RGBD, Pose 125000 50 500 CSL - Lab
@@ 8.2 Semaphoric HGR
Semaphoric HGR research is supported by a diverse range of datasets categorized into three main types based on task formulation: SHGR, IDHGR and CDHGR. These datasets vary in their application domains (e.g., HCI, HMI, HRI), modalities (e.g., RGB, depth, pose, infrared), and in the number of gesture classes, subjects, and gesture samples they include. The statistics of datasets for semaphoric HGR are provided in tables 7,8 and 9, regarding the task for which they are used (i.e. static, isolated dynamic and continuous HGR, respectively).
SHGR datasets target gesture recognition from single frames, making them easier to process due to the lack of temporal complexity. Older datasets like HGR-1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands) contain few samples, limiting deep learning training, yet remain benchmarks for HCI due to their simplicity. LaRED (LaRED) offers 243K RGBD samples across 81 gestures, but only from 10 subjects, restricting subject-independent recognition. HaGRID (WACV_kapitanov2024hagrid) is much larger and diverse, with nearly twice as many full-resolution RGB frames from  35K subjects in varied scenes, including bounding boxes for hand detection. HaGRIDv2 (HAGRIDv2_ARXIV) expands this with 15 additional classes,  66K subjects, and over 1M samples. SHAPE (SHAPE_DATASET) has a similar number of gestures (all two-handed) but fewer samples. LRHG (LONG_RANGE_HGR_SSD) and MD-UHGRD (MD-UHGRD_SA-YOLO) target UAV-based gesture recognition, with MD-UHGRD offering more samples but fewer classes, captured from varying distances and angles to enhance viewpoint robustness. URGR (ULTRA-RANGE-UGV-Bamani) focuses on ultra-range gestures for UGVs, with 6 robot-command gestures performed by 16 signers at distances from 2 to 25 meters.
IDHGR datasets contain gestures performed in isolation, with start and end points explicitly segmented. Most of the existing datasets (e.g. SKIG (SKIG), NVIDIA (aka nvGesture) (NVIDIA_GESTURE_R3DCNN_Molchanov), DHG (DHG-CVPRW) and SHREC17 (SHREC)) for IDHGR are relatively small with a few thousands samples and support multimodal settings by providing depth maps, IR frames, optical flow or skeleton joints, accompanying the RGB modality aiming to enhance the accuracy of the models. Jester-20bn (Jester-Dataset), proposed for gesture-based HCI, is the largest dataset for (non-SLU) IDHGR containing 1̃50K videos performed by 1376 subjects. The remarkable diversity and size of this dataset make it ideal for pretraining DL models. Other datasets for HCI include IPN-Hand (IPN-HAND), ZJUGesture (ZJUGesture) and DATE (DATE), comprising 4218, 9892 and 13500 samples respectively. LD-ConGR (LD-ConGR), ChaLearn IsoGD (chalearn-lap) and EgoGesture (EgoGesture_DATASET) have sufficient amount of data, constituting adequate baseline for IDHGR. EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET), Briareo (Briareo_dataset) and nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) have been proposed for human-car interaction, mainly in multimodal setting, including gestures that correspond to car functionalities (e.g. CW rotation and audio manipulation [^2] https://youtu.be/NJmk1DUyyB8). TCG (TCG) was also proposed for human-car interaction, but for recognizing traffic police gestures in autonomous driving scenarios relying only on pose data. EHWGesture (EHWGesture-amprimo2025ehwgesture) is the first dataset for recognizing clinical gestures and providing annotations for many other tasks, including gesture detection & triggering.
CDHGR datasets challenge models to identify and segment gestures from continuous video streams. Many of the datasets mentioned earlier (e.g. ChaLearn (chalearn-lap), IPN-Hand (IPN-HAND), LD-ConGR (LD-ConGR), EgoGesture (EgoGesture_DATASET) and ZJUGesture (ZJUGesture)) have been obtained by annotating the temporal boundaries of gestures within continuous sequences, thus these datasets have protocols for both IDHGR and CDHGR. Therefore, we will only discuss the Police Gesture Dataset (Police-Gesture-Dataset) as the previous analysis could also be applied here. This dataset was introduced for recognizing police traffic gestures, similarly with TCG (TCG), containing a small number of samples (21 RGB videos) and 3354 gesture instances.
@@
Table 7: Datasets for SHGR.
Name Year Domain Modality Samples Subjects Classes Resolution Distance
HaGRID (WACV_kapitanov2024hagrid) 2024 HCI RGB 552992 34730 19 1920x1080 0.5-4m
HaGRIDv2 (HAGRIDv2_ARXIV) 2024 HCI RGB 1086158 65977 34 1920x1080 0.5-4m
MD-UHGRD (MD-UHGRD_SA-YOLO) 2024 HRI RGB 20000 - 19 - 3-10m
URGR (ULTRA-RANGE-UGV-Bamani) 2024 HRI RGB 347483 16 6 640x480 1-25m
SHAPE (SHAPE_DATASET) 2022 - RGB 34226 20 32 4128x3096 2-5m
LRHG (LONG_RANGE_HGR_SSD) 2021 HRI RGB 4320 8 10 640x480 1-7m
OUHANDS (ouhands) 2016 HCI RGBD 3150 23 10 640x480 $\leq$ 1m
HGR-1 (HGR-1-2024-PATTER-RECOGNITION) 2014 - RGB 899 12 27 174x131-640x480 $\leq$ 1m
LaRED (LaRED) 2014 - RGBD 243000 10 81 640x480 -
NUS-II (NUS-II-2013) 2013 - RGB 2000 40 10 160x120 $\leq$ 1m
ASL-FS (ASL-FS-2011-CVPRW) 2011 - RGBD 48K 4 24 - -
NUS-I (NUS-I-2010) 2010 - RGB 240 - 10 160x120 $\leq$ 1m
Table 8: Datasets for IDHGR.
@@
EHWGesture (EHWGesture-amprimo2025ehwgesture) 2025 - RGBD 1100 25 11 3.3M
EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET) 2024 HMI RGB 400 - 29 70276
DATE (DATE) 2024 HCI RGB 13500 22 27 -
ZJUGesture (isol.) (ZJUGesture) 2023 HCI RGB 9892 60 9 -
LD-ConGR (isol.) (LD-ConGR) 2022 HCI RGBD 44887 30 10 -
IPN Hand (isol.) (IPN-HAND) 2021 HCI RGB 4218 50 13 -
TCG (TCG) 2020 HMI Pose 250 5 4 839350
Jester (Jester-Dataset) 2019 HCI RGB 148092 1376 27 5331312
Briareo (Briareo_dataset) 2019 HMI RGBD, Pose, IR 1440 40 12 -
EgoGesture (isol.) (EgoGesture_DATASET) 2018 HCI RGBD 24161 50 83 2953224
SHREC17-14/28 (SHREC) 2017 - Depth, Pose 2800 28 14/28 -
nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) 2016 HMI RGBD, IR, OF 1532 20 25 -
ChaLearn IsoGD (chalearn-lap) 2016 - RGBD 47933 21 249 -
DHG-14/28 (DHG-CVPRW) 2016 - Depth, Pose 2800 20 14/28 -
SKIG (SKIG) 2013 - RGBD 1080 6 10 -
Table 9: Datasets for CHGR.
@@
ZJUGesture (cont.) (ZJUGesture) 2023 HCI RGB - 9892 60 9
LD-ConGR (cont.) (LD-ConGR) 2022 HCI RGBD 542 44887 30 10
IPN Hand (cont.) (IPN-HAND) 2021 HCI RGB 200 4218 50 13
Police Gesture Dataset (Police-Gesture-Dataset) 2020 HMI RGB 21 3354 - 8
EgoGesture (cont.) (EgoGesture_DATASET) 2018 HCI RGBD 2081 24161 50 83
ChaLearn ConGD (chalearn-lap) 2016 - RGBD 22535 47933 21 249
@@ 8.3 Other
In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types. The First-Person Hand Action (FPHA) dataset (FPHA), published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants. It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction. Another notable example is the DP Dataset (DeePoint), released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition. Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances. It includes detailed annotations such as gesture timings and pointing directions, enabling both binary gesture classification (gesture vs. non-gesture) and directional estimation tasks. Finally, the SocialGesture (SocialGesture-CVPR2025-cao2025socialgesture) focuses on the recognition of natural gestures in multi-person environments. It comprises 9,889 in-the-wild videos containing 42,533 instances and provides fine-grained annotations including bounding boxes, relations (for example, a subject is pointing to a target) and Visual Question Answering (VQA) details.
8.4 Datasets Comparison
After examining the existing datasets for VHGR with significant impact on the current research, in this subsection we aim to determine pros and cons of the most important ones as well as the use cases that they fit. For gesture-based HCI, HaGRID (WACV_kapitanov2024hagrid) and its versions (HAGRIDv2_ARXIV) seem to be a reliable solution, containing an enormous number of samples corresponding to 19 distinct classes (that can be mapped to computer commands). Regarding that these gestures are also static, even simple architectures - like the ones discussed in sections 4 & 7 - can achieve sufficient performance. Furthermore, this dataset could also be used for pretraining, enhancing the backbone’s expression capability. However, its size implies that training takes a long time, which constitutes a deterrent in some cases, compared to other datasets for HCI like OUHANDS (ouhands). For gesture-based HRI, URGR (ULTRA-RANGE-UGV-Bamani) is the only one for UGV guidance and considers various recognition ranges, which may be crucial for real-world scenarios. Even though it contains a large number of samples, the small number of classes may limit the commands given by gestures and, therefore, the functionalities of the entire gesture-based interface. Regarding SLU datasets, for ASL recognition, OpenASL (OpenASL) is the largest one, including almost 100K RGB videos (288h in total) collecting from the web. MSASL (MS-ASL) comprises almost the same number of subjects, but it is much smaller. Furthermore, OpenASL was proposed for CSLR while for MSASL for IDSLR, thus it may be more suitable for real-world use cases, where the temporal boundaries are not known. BSL-1K (bsl-1K) is also a very large dataset, proposed for British IDSLR and comprising the largest number of samples (280K, 1060h in total). BOBSL (BOBSL-BBC) was proposed for British CSLR and comprises videos of 1467h in total making it the dataset with the longest duration. Phoenix (PHOENIX-2014) and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) remain the most adequate dataset for German sign language, covering a large sign vocabulary. Nevertheless, the small number of samples and the low diversity - due to the limited number of signers and the constrained environment - make essential the collection of a new, larger and more diverse dataset for DGS. For Chinese SLR, NMFs-CSL (NMFs-CSL) is significantly smaller than CSLD (tim-slr_csld_dataseta), CSL (CSL500) and NCSL (2plus1-D-SLR_and_NCSL) but includes much more signs (1064 instead of less than 500). CSL-Daily (Back-Translation-CVPR) remains the only one for Chinese CSLR, covering a large sign vocabulary of 2000 signs in total.